### v4

# Testing LRD in the spectral domain for functional time series in manifolds

Abstract
A statistical hypothesis test for long range dependence (LRD) is formulated in the spectral domain for functional time series in manifolds. The elements of the spectral density operator family are assumed to be invariant with respect to the group of isometries of the manifold. The proposed test statistic is based on the weighted periodogram operator. A Central Limit Theorem is derived to obtain the asymptotic Gaussian distribution of the proposed test statistic operator under the null hypothesis. The rate of convergence to zero, in the Hilbert–Schmidt operator norm, of the bias of the integrated empirical second and fourth order cumulant spectral density operators is obtained under the alternative hypothesis. The consistency of the test follows from the consistency of the integrated weighted periodogram operator under LRD. Practical implementation of our testing approach is based on the random projection methodology. A simulation study illustrates, in the context of spherical functional time series, the asymptotic normality of the test statistic under the null hypothesis, and its consistency under the alternative. The empirical size and power properties are also computed for different functional sample sizes, and under different scenarios.
M.D. Ruiz-Medina1 and Rosa M. Crujeiras2
1 University of Granada
2 University of Santiago de Compostela
MSC2020 subject classifications:
Primary. 60G10, 60G12, 60G18 , 60G20, 60G22
Secondary. 60G60
Keywords. Asymptotic normality; bias; compact manifolds; consistency; empirical cumulant spectral density operator; functional time series; integrated weighted periodogram operator; long-range dependence; spectral density operator.
1 Introduction
Spherical functional time series analysis helps in understanding the dynamics and spatiotemporal patterns of data that are embedded into the sphere, providing valuable insights for prediction, monitoring, and decision-making. Time series analysis of global temperature data distribution among other climate variables, usually arising in Climate Science and Meteorology, can be performed in a more efficient way by adopting a functional time series framework (see [38]). That is the case of ocean currents, and other marine functional time series to be analyzed in Oceanography studies (see, e.g., [41]; [43]). Other areas demanding this type of techniques are Geophysics, Astronomy and Astrophysics. In the last few decades, the cosmic microwave background radiation variation analysis over time has gained special attention (see [13]; [22]; [23]). In a more general manifold setting, functional time series analysis is often applied in Medical Imaging, Computer Vision and Graphics (see [42]; [45]; [44], among others). This paper focuses on the spectral analysis of functional time series in manifolds, with special attention to LRD analysis.
The spectral analysis of functional time series has mainly been developed under Short Range Dependence (SRD). In this context, based on the weighted periodogram operator, a nonparametric framework is adopted in [28]. Particularly, the asymptotic normality of the functional discrete Fourier transform (fDFT), and the weighted periodogram operator of the curve data are proved under suitable summability conditions on the norm of the cumulant spectral density operators. The consistency of the weighted periodogram operator, in the Hilbert–Schmidt operator norm, is derived under SRD. In [29], a harmonic principal component analysis of functional time series in the temporal functional spectral domain is also obtained, based on a Karhunen–Loéve–like decomposition, the so–called Cramér–Karhunen–Loéve representation. In the context of functional regression, some applications are presented in [30], [34] and [39]. Hypothesis testing for detecting modelling differences in functional time series dynamics is addressed in [40] in the functional spectral domain.
In LRD analysis of functional time series several problems still remain open. One of the key approaches in the current literature is presented in [19], where the eigendecomposition of the long-run covariance operator is considered, under an asymptotic semiparametric functional principal component framework. The consistent estimation of the dimension and the orthonormal functions spanning the dominant subspace, where the projected curve process displays the largest dependence range is derived. Fractionally integrated functional autoregressive moving averages processes constitute an interesting example of this modelling framework.
A first attempt to characterize LRD in functional time series in the spectral domain can be found in [36], adopting the theoretical framework of operator–valued random fields, including fractional Brownian motion with operator–valued Hurst coefficient (see, e.g., [12]; [24]; [31] and [32]). The eigenvalues of the LRD operator are parameterized. These eigenvalues induce different levels of singularity at zero frequency, corresponding to different levels of temporal persistent of the process projected into different eigenspaces of the Laplace Beltrami operator. Under this LRD scenario, the integrated periodogram operator is proved to be asymptotically unbiased in the Hilbert–Schmidt operator norm. Minimum contrast estimation of the LRD operator is achieved in the spectral domain in a weak–consistent way under a Gaussian scenario. Interesting examples of this setting are analyzed in [26], where the spectral analysis of multifractionally integrated functional time series in manifolds is considered. In particular, multifractionally integrated spherical functional ARMA models (i.e., multifractionally integrated SPHARMA models) are analyzed through simulations. In this modelling framework, SRD and LRD can coexist at different spherical scales.
Up to our knowledge, no further developments have been achieved in the spectral analysis of LRD functional time series. Alternative contributions for stationary LRD functional sequences are based on the diagonalization of the heavy tail autocovariance kernel of the time–varying functional error term (see, e.g., [3]). However, under this modelling framework, functional spectral analysis can not be achieved in the time domain due to the assumed independence between the random components of the error term (see also [2]). Similar assertions hold for the two sample problem analyzed in [4]. In [20], a special family of LRD linear functional time series is analyzed with scalar coefficients displaying slow decay. Under stationarity, this LRD scenario constitutes a particular case of our framework when the elements of the spectral density operator family have degenerated pure point spectra corresponding to one infinite–dimensional eigenspace. The Cramér–Karhunen–Loéve representation above referred constitutes a powerful tool in the functional spectral analysis of weak-dependent functional time series (see [35]). In this paper, this representation is extended to our LRD stationary functional time series context, assuming the invariance of the cross-covariance kernels under the group of isometries of the manifold, given by a connected and compact two points homogeneous space.
In this paper we perform a weighted periodogram operator based analysis, requiring the asymptotic analysis of the bias of the integrated empirical fourth–order cumulant spectral density operators, to prove consistency of the integrated weighted periodogram operator under LRD. Its application to spectral statistical hypothesis testing of LRD in –valued correlated sequences constitutes one of the main goals of this work. Here, denotes the space of real–valued square integrable functions on a Riemannian manifold embedded into given by a connected and compact two–point homogeneous space. The topological dimension of is and denotes the normalized Riemannian measure on In what follows, we will consider to be a functional sequence such that for every with denoting the probability measure defined on the basic probability space i.e., for every
is a measurable mapping.
The invariance of the elements of the spectral density operator family of under the group of isometries of the manifold is assumed along the paper. A frequency–varying eigenvalue sequence then characterizes the pure point spectra of the elements of the spectral density operator family, with respect to the orthonormal basis of eigenfunctions of the Laplace–Beltrami operator. This invariance assumption is exploited in the derived Central Limit Theorem that characterizes the asymptotic distribution of the proposed test statistic operator under the null hypothesis, which states that displays SRD. In our formulation of the alternative hypothesis on LRD, we adopt a semiparametric framework in terms of a functional parameter given by the LRD operator. In contrast with the approach presented in [36], here we do not assume a parameterization of the eigenvalues of the LRD operator. Under this scenario, the rate of convergence to zero, in the corresponding norm, of the the bias of the integrated empirical second and fourth order cumulant spectral density operators is respectively obtained in Lemmas 2 and 3 under suitable conditions. Proposition 1 shows the divergence, in the Hilbert–Schmidt operator norm, of the mean of the test statistic operator under the alternative hypothesis. Theorem 2 derives suitable conditions, in particular, on the nonparametric functional spectral factor, to ensure consistency of the integrated weighted periodogram operator in the Hilbert–Schmidt operator norm under LRD. Theorem 3 then provides the almost surely divergence of the test statistic in the Hilbert–Schmidt operator norm under the alternative, yielding the consistency of the test.
Theorem 3 also plays a crucial role in the implementation in practice of the proposed testing procedure, based on rejecting the null hypothesis when the random Fourier coefficients of the test operator statistic, suitably standardized according to Theorem 1, cross an upper or lower tail standard normal critical value. The orthonormal basis involved in the computation of these coefficients is constructed by the tensor product of the eigenfunctions of the Laplace Beltrami operator. The random projection methodology (see Theorem 4.1 in [8]) can be implemented here to alleviate the dimensionality problem. Specifically, when the moments of our test statistic operator satisfy the Carleman condition under the null hypothesis, our testing procedure is equivalent to rejecting the null hypothesis when the absolute value of a random projection of the test statistic is larger than an upper tail standard normal critical value. In the implementation of the random projection methodology, the Karhunen–Loéve expansion in Lemma 4 below can be considered for generation of the involved Gaussian random directions. In the simulation study undertaken, robust empirical sizes, and competitive values of the empirical power are displayed by our testing approach (see Section 5.4).
The outline of the paper is as follows. In Section 1.1, the functional spectral background material is introduced. Our hypothesis testing procedure is formulated in a functional semiparametric spectral framework in Section 1.2. The asymptotic Gaussian distribution of the test statistic operator under the null hypothesis is obtained in Theorem 1 in Section 2. Asymptotics of the bias of the integrated empirical second and fourth order cumulant spectral density operators under LRD are derived in Section 3. Section 4 provides the preliminary results required for consistency of the test, which is derived in Theorem 3 of this section. Practical implementation is also discussed in Section 4. In Section 5.1, a simulation study is undertaken to illustrate the asymptotic Gaussian distribution of the proposed test statistic operator under the null hypothesis, in the context of SRD spherical functional time series. The consistency of the test is also illustrated in Section 5.2, in the framework of multifractionally integrated spherical functional time series. This numerical analysis is extended in Section 5.3 to a wider family of LRD operators allowing stronger persistency in time, displayed by the projected process in the dominant subspace. Section 5.4 analyzes empirical size and power properties of the test. Section 6 summarizes conclusions of the simulation study, focusing on the large functional sample size properties of our test statistic operator under different bandwidth parameter scenarios, from additional numerical results. The proofs of the results of this paper can be found in the Appendix.
1.1 Background
Along this work we will assume that in (1) is a stationary zero–mean functional sequence, with nuclear covariance operator family satisfying for every The elements of this family are characterized in the spectral domain by the spectral density operator family The assumed invariance of the elements of these families with respect to the group of isometries of yields to their diagonal series expansion in terms of with being the orthonormal basis of eigenfunctions of the Laplace–Beltrami operator on (see, e.g., [15]; [17]). In particular,
where, for every denotes the dimension of the eigenspace associated with the eigenvalue of the Laplace Beltrami operator (see, e.g., Section 2.1 in [21]). The equality means identity in the norm of the space of Hilbert–Schmidt operators on the space of complex–valued square integrable functions on Specifically, the equality in (2) means that
where is the kernel of the integral operator for every
Let be a functional sample of size of The fDFT is defined as
The kernel of the periodogram operator satisfies, for each
We will denote by the kernel of the cumulant operator of order of the fDFT over the diagonal Note that, for and the Féjer kernel is given by
The weighted periodogram operator, denoted as has kernel given, for each by
where is a weight function satisfying
with being the positive bandwidth parameter. Function is a real–valued function defined on such that is positive, even, and bounded in variation; , if ;
1.2 Hypothesis testing
The SRD and LRD scenarios respectively tested under the null and the alternative hypotheses are introduced in this section. The proposed test statistic operator based on the weighted periodogram operator is then formulated.
Stationary SRD functional time series are characterized by the summability of the series of trace norms of the elements of the family of covariance operators (see, e.g., [28]). That is, displays SRD if and only if where denotes the space of trace operators on In our setting, this condition can be formulated as follows:
When (8) fails, is said to display LRD. In what follows, we will adopt the LRD scenario introduced in [36], given by
where the invariant positive self–adjoint operators and are composed yielding the definition of Specifically, denotes the LRD operator on Operator in (9) is interpreted as in [7], [31] and [32], where plays the role of operator–valued Hurst coefficient in the setting of fractional Brownian motion introduced in this framework. Moreover, is the regular spectral operator reflecting markovianess when the null space of coincides with Specifically, satisfies
where the operator integrals are understood as improper operator Stieltjes integrals which converge strongly (see, e.g., Section 8.2.1 in [33]).
We will apply the spectral theory of self–adjoint operators (see, e.g., [11]) in terms of the common spectral kernel
under the assumed invariance property with respect to the group of isometries of
The point spectrum of is given by with for every and It is assumed that LRD operator has kernel admitting the following series expansion in the weak–sense:
Specifically, identity (11) is understood as
for all where denotes the space of infinitely differentiable functions with compact support contained in Note that, under the conditions assumed, and are bounded, and with denoting the norm in the space of bounded linear operators on
In a similar way, operator is interpreted as
for every and
Operator in (9) is a Hilbert–Schmidt operator on whose kernel admits the following series expansion in the norm of the space
where denotes the sequence of positive eigenvalues. For each is a continuous positive slowly varying function at in the Zygmund’s sense (see Definition 6.6 in [1], and Assumption IV in [36]). Equation (10) can be equivalently expressed, in terms of as
As before, equation (15) implies that displays SRD, when for every Under (15), is also included in the trace class.
Under the above setting of conditions,
i.e., with being the space of square integrable functions on the interval Condition (16) plays a crucial role in the derivation of the results of this paper under LRD.
From equations (9)–(14), the elements of the positive frequency varying eigenvalue sequence in (2) admit the following expression:
Note that, since
Sequence (17) is involved in the formulation of the alternative hypothesis stating that displays LRD against where SRD is assumed. Specifically,
In our context, the formulation of the test statistic must capture the singularities at zero frequency for different manifold resolution levels under The proposed test statistic operator is then given by
where the kernel of the integral operator has been introduced in equation (6), with, as before, being the bandwidth parameter. The indicator function on the interval is denoted by for Note that, as
for every (see [14] for the usual notion of convergence in the sense of generalized functions or distributions). Here, denotes the Dirac Delta distribution at zero frequency. Hence, in what follows we adopt the notation representing a truncated Dirac Delta distribution.
2 Preliminary results under SRD
The following lemma will be applied in the proof of the main result of this section, Theorem 1, deriving the asymptotic Gaussian distribution of in (21) under Specifically, Lemma 1 provides the asymptotic Gaussian distribution of the weighted periodogram operator under Its proof can be obtained in the same way as in [28], where this result is formulated for the separable Hilbert space
Lemma 1
Assume that for all and
(i′)
for -
Then, for every frequencies with
where denotes the convergence in distribution. Here, are jointly zero–mean complex Gaussian elements in with covariance kernel:
with for and otherwise. Thus, and are independent for mod and mod For zero frequency modulus the limit Gaussian random element is in
Proof. See Theorem 3.7 in [28].
The next result provides the asymptotic Gaussian distribution of the test statistic operator under The convergence to a Gaussian random element in the norm of the space is also obtained. Here, denotes the space of zero–mean second–order –valued random variables with the norm
Theorem 1
Proof. See Appendix A.1.
3 Second and fourth order bias asymptotics under LRD
This section provides new results on the bias asymptotics in the Hilbert-Schmidt operator norm of the integrated empirical second and fourth order cumulant spectral density operators of under These results are applied in the derivation of Theorem 2 and Corollary 2, providing the consistency of the integrated weighted periodogram operator under In what follows, we assume and as
The rate of convergence to zero of the norm of the bias in the space of the integrated periodogram operator is obtained under LRD in the next lemma. The following well-known identity will be applied:
for where denotes the Féjer kernel introduced in equation (5) of Section 1.1.
Lemma 2
Under as
where denotes expectation under the alternative and, as before, denotes the equality in the norm of the space
Proof. See Appendix B.1.
The following corollary is obtained from Lemma 2, and provides the rate of convergence to zero of the bias of the integrated weighted periodogram operator, in the norm of the space under
Corollary 1
Under as
Proof. See Appendix B.2.
The rate of convergence to zero, in the norm of the space of the bias of the integrated empirical fourth–order cumulant spectral density operators of under LRD is obtained in Lemma 3 below. The following assumption is required:
Assumption I. For every defines an isotropic kernel in and the following convergence holds:
where
Lemma 3
Under and Assumption I, uniformly in
where, for
denotes the cumulant spectral density operator of order of and, as before, means the identity in the norm of the space
Proof. See Appendix B.3.
4 A test for LRD in functional time series on
Consistency of the test based on is derived in this section. Specifically, Theorem 3 provides the almost surely divergence in the norm of the space of under The proof of this result follows from Proposition 1, showing the divergence in the norm of the space of the centering operator of and from Theorem 2 and Corollary 2, establishing the consistency of the integrated weighted periodogram operator under The implementation of the testing procedure in practice is also discussed.
Proposition 1
Under as
Proof. See Appendix C.1.
Theorem 2
Under Assumption I, and
as
where, as before, denotes the norm in the space of nuclear operators on
Remark 1
Note that condition (32) is satisfied, for instance, when the family lies in a ball of radius of the space
Proof. See Appendix C.2.
Theorem 2 implies the weak consistency of the integrated weighted periodogram operator under in the norm of the space
Corollary 2
Under the conditions of Theorem 2, as
Proof. See Appendix C.3.
Under the conditions assumed in the following result, consistency of the test follows.
Theorem 3
Under assume that and that the bandwidth parameter for If conditions of Theorem 2 hold, then, as
where denotes almost surely divergence.
Proof. See Appendix C.4.
4.1 Practical implementation
The practical implementation of the proposed statistical testing procedure, in terms of the Gaussian random projection methodology (see Theorem 4.1 in [8]), is now briefly discussed. The Karhunen–Loéve expansion in Lemma 4 below can be applied in such an implementation. Let us consider the random Fourier coefficients
where integration is understood in the mean–square sense, and is the limit Gaussian random element in introduced in Lemma 1.
Lemma 4
Let be defined, as before, satisfying equation (23) in Lemma 1. Then, the following series expansion holds in the mean–square sense: For every
where, as before, denotes the space of zero–mean second–order –valued random variables with the norm The random Fourier coefficients
for have been introduced in equation (34). They are independent and identically distributed complex–valued standard Gaussian random variables.
Proof. See Appendix C.5.
Theorem 3 motivates the methodology to be adopted in practice. Specifically, as illustrated in the simulation study undertaken in the next section, a consistent test for LRD is obtained by rejecting when, for every
is larger than an upper tail standard normal critical value. Note that for sufficiently large, and for
The associated dimensionality problem can be substantially alleviated if we restrict our attention to the case where all moments of are finite and satisfy the Carleman condition. In that case, Theorem 4.1 in [8] leads to the following test statistic, evaluated conditionally to the observed functional value of a non–degenerated functional Gaussian random variable in the space whose probability measure on is denoted as Specifically, consider
Then, will be rejected if the observed value of is larger than an upper tail standard normal critical value. Note that, if holds then also holds, and if fails then also fails –a.s. Thus, with probability one, we will generate a realization of random direction in for which fails (see also [9]).
In the spirit of the Gaussian random degree– spherical harmonics introduced in [25] (), we will consider a truncated Karhunen–Loéve expansion in the generation of a non–degenerated Gaussian measure characterizing the random direction where our test statistics is projected (see equation (38)). Specifically, we consider a zero–mean Gaussian random element in with trace covariance operator having kernel
for every The trace property of can be equivalently expressed as Therefore, admits a series expansion, Karhunen–Loéve expansion (see, e.g., Lemma 4), whose truncated version is implemented. Particularly, one can construct centered isotropic Gaussian random fields by finite–dimensional projection
for and involving the Gaussian random directions in
where denotes a zero–mean Gaussian random variable with variance for Kernel plays a key role in our approach when and/or are dominant eigenspaces of the Laplace Beltrami operator on
5 Simulation study
Our simulations will be set on An alternative generation algorithm to the ones considered in [26] and [27] is implemented reducing computational burden allowing for the consideration of large functional sample sizes. Theorem 1 is illustrated in the context of SPHARMA(p,q) processes, and the illustration of Theorem 3 is carried out in the context of multifractionally integrated SPHARMA(p,q) processes. These numerical results are respectively reported in Sections 5.1 and 5.2 for i.e., (see also Section 6 for a wider analysis for different values). When Section 5.3 opens new research lines beyond condition (16), providing empirical evidence of a faster a.s. divergence rate of in the norm of the space Finally, Section 5.4 shows empirical size and power properties of the testing procedure.
5.1 Asymptotic Gaussian distribution of under
Let us consider that the elements of the family of spectral density operators of have frequency varying eigenvalues, with respect to the system of eigenfunctions of the Laplace–Beltrami operator, obeying the following equation under (see [36]):
where is the system of eigenvalues of the autocovariance operator of the innovation process with respect to the system of eigenfunctions of the Laplace–Beltrami operator. Process is assumed to be strong–white noise in That is, is assumed to be a sequence of independent and identically distributed –valued random variables such that and with and for and for For and with and denoting the sequences of eigenvalues, with respect to the system of eigenfunctions of the Laplace–Beltrami operator, of the self–adjoint invariant integral operators and for and respectively. These operators satisfy the following equations:
where is a difference operator such that
Here, and are the so–called autoregressive and moving average operators, respectively. Also, for each and have not common roots, and their roots are outside of the unit circle (see also Corollary 6.17 in [1]). Thus, satisfies an SPHARMA(p,q) equation (see also [5]; [6]).
In the simulations we have generated an SPHARMA(1,1) process, i.e., with and and Figure 1 displays one realization of the generated SPHARMA(1,1) process projected into at times
For each the empirical distribution of the centered and standardized projection into of is displayed in Figure 2 for functional sample size and repetitions, in Figure 3 for functional sample size and repetitions, and in Figure 4 for functional sample size and repetitions. These empirical distributions approximate the support and shape of a standard Gaussian probability density. The empirical standardization displays a decreasing pattern over the spherical scale meaning that the respective limit one–dimensional Gaussian probability measures of these projections have decreasing support. According to Theorem 1.2.1 in [10], this property is satisfied by the infinite product Gaussian measure on whose restriction to is identified in the –sense with the probability measure of the limit Gaussian random element in Theorem 1.
5.2 Consistency of the test
Under for each the eigenvalues satisfy (see equations (17), (18) and (20))
Again, we consider the projection of into Three multifractional integration operators, applied to SPHARMA(1,1) process generated in Section 5.1, are considered in Sections 5.2.1–5.2.3. In Example 1, is decreasing over in Example 2 is increasing over and non–monotone in Example 3. Note that, under the generated Gaussian scenario, condition (16) implies that Assumption I holds. Furthermore, condition (32) also holds since lies in the unit ball of the space and belongs to the trace class.
5.2.1 Example 1
Theorem 3 is now illustrated in the case where the largest dependence range is displayed by the process projected into the eigenspace Figure 5 displays a sample realization of the corresponding multifractionally integrated SPHARMA (1,1) process projected into
In this example, and (see plot at the left–hand side of Figure 6).
5.2.2 Example 2
The dominant subspace in this example, where the projected process displays the largest dependence range, is eigenspace One sample realization of the generated multifractionally integrated SPHARMA (1,1) process, projected into is displayed in Figure 7.
The LRD operator eigenvalues are given in the plot at the left–hand side of Figure 8, where and The a.s. divergence of our test statistic operator in the Hilbert–Schmidt operator norm (see also Table 1) is illustrated in the three plots at the right–hand side of such Figure 8, in terms of the sample values of each projection of into for increasing functional samples sizes
5.2.3 Example 3
In this third example, the dominant subspace is eigenspace of the Laplace–Beltrami operator One sample realization of the generated multifractionally integrated SPHARMA (1,1) process, projected into is displayed in Figure 9.
The eigenvalues of LRD operator are showed in the plot at the left–hand side of Figure 10 with and The sample values of the projections of into for functional samples sizes can be found in the three plots at the right–hand side of Figure 10 (see also Table 1).
5.2.4 Almost surely divergence of in norm under
The observed values of the Hilbert-Schmidt operator norm of for projected into is displayed in Table 1, for the three numerical examples generated, and for the functional sample sizes
| T | Example 1 | Example 2 | Example 3 |
| 1000 | 2.3036e+05 | 2.2595e+05 | 1.9934e+05 |
| 5000 | 1.0612e+07 | 1.0223e+07 | 9.0697e+06 |
| 10000 | 5.5172e+07 | 5.3770e+07 | 4.7303e+07 |
| 30000 | 7.5695e+08 | 7.3377e+08 | 6.4742e+08 |
| 50000 | 2.5516e+09 | 2.4764e+09 | 2.1844e+09 |
| 100000 | 1.3256e+10 | 1.2892e+10 | 1.2906e+10 |
One can observe in Table 1 the increasing sample values of the Hilbert–Schmidt operator norm of the projected as increases, in all the examples under with satisfying Spherical sample patterns and scales induced by the multifractional integration operator (see Figures 5, 7 and 9) have no significant effect (see Table 1), when the condition is satisfied under the bandwidth parameter modelling This fact is also reflected in Figures 6, 8 and 10, respectively, where decreasing patterns, and almost the same divergence rates are displayed by the sample values of projected into for in all the examples. However, the scenario under which crosses the threshold at some spherical scale requires a separated analysis, as briefly discussed in Example 4 in the next section (see Figure 12).
5.3 Example 4
Our numerical analysis is extended here beyond the restriction Specifically, this section shows some preliminary numerical results regarding the effect of higher levels of singularity at zero frequency when i.e., corresponding to a stronger persistency in time of the projected process into the dominant subspace (see Figure 11).
Under this scenario, conditions (16), and (32) in Theorem 2, are not satisfied. Indeed, we are out of the scenario where the summability in time of the square of the Hilbert–Schmidt operator norms of the elements of the covariance operator family holds. Then, new technical tools are required to address the asymptotic analysis in the spectral domain of this family of manifold supported functional time series displaying stronger levels of persistency in time. Let us again consider for projected into In this example, the multifractional integration of SPHARMA(1,1) process generated in Section 5.1 has been achieved in terms of LRD operator having eigenvalues displayed at the left–hand side of Figure 12, with and and being the dominant subspace. The same functional sample sizes as in Examples 1–3 have been considered. One can observe, in the three plots displayed at the right–hand side of Figure 12 , that the decreasing patterns over displayed in Figures 6, 8 and 10 do not hold in this example. Table 2 also illustrates a faster increasing than in Examples 1–3 of for functional sample sizes ranging from to under and
( and )
| Sample | |||||
| Size | |||||
| 1000 | 5000 | 10000 | 30000 | 50000 | 100000 |
| 6.5651e+05 | 3.8623e+07 | 2.2172e+08 | 3.5383e+09 | 1.2688e+10 | 7.2258e+10 |
5.4 Empirical size and power analysis
The empirical size and power properties of the testing approach presented are now illustrated. We have applied the random projection methodology. Tables 3 and 4 display the numerical results for random functional directions (see equation (38)). Model SPHARMA(1,1) generated in Section 5.1 has been considered in the computation of the empirical size of the test. Multifractionally integrated SPHARMA(1,1) model, generated in Section 5.2.1, defines the scenario under the alternative to compute the empirical power. For each one of the eight random directions tested, we have analyzed the functional samples sizes and, for each functional sample size, we have considered repetitions.
The empirical size properties of the proposed testing procedure are quite robust, as one can observe in the numerical results displayed in Table 3. Specifically, since we are working with finite sample sizes (), despite the statistical distance to the normal distribution which holds asymptotically, small deviations are observed from the theoretical value for all number of repetitions considered. One can also observe, in Table 4, the increasing patterns displayed by the empirical power with respect to the functional sample sizes tested in all random directions. Note that these empirical power values are in the interval In particular, since the threshold the empirical power is almost for any of the three values of studied.
| R | T 50 | |||||||
| 500 | 0.0280 | 0.0560 | 0.0360 | 0.0480 | 0.0600 | 0.0600 | 0.0360 | 0.0520 |
| 1000 | 0.0480 | 0.0420 | 0.0320 | 0.0380 | 0.0420 | 0.0440 | 0.0320 | 0.0300 |
| 3000 | 0.0420 | 0.0447 | 0.0453 | 0.0353 | 0.0413 | 0.0400 | 0.0440 | 0.0507 |
| R | T 100 | |||||||
| 500 | 0.0360 | 0.0680 | 0.0440 | 0.0720 | 0.0520 | 0.0240 | 0.0360 | 0.0400 |
| 1000 | 0.0280 | 0.0380 | 0.0380 | 0.0500 | 0.0380 | 0.0740 | 0.0340 | 0.0360 |
| 3000 | 0.0373 | 0.0507 | 0.0407 | 0.0460 | 0.0440 | 0.0360 | 0.0600 | 0.0480 |
| R | T500 | |||||||
| 500 | 0.0440 | 0.0520 | 0.0320 | 0.0640 | 0.0320 | 0.0480 | 0.0320 | 0.0480 |
| 1000 | 0.0420 | 0.0400 | 0.0500 | 0.0460 | 0.0420 | 0.0380 | 0.0540 | 0.0240 |
| 3000 | 0.0453 | 0.0393 | 0.0447 | 0.0407 | 0.0427 | 0.0507 | 0.0453 | 0.0553 |
| R | T1000 | |||||||
| 500 | 0.0520 | 0.0360 | 0.0400 | 0.0560 | 0.0600 | 0.0640 | 0.0480 | 0.0520 |
| 1000 | 0.0440 | 0.0380 | 0.0400 | 0.0580 | 0.0500 | 0.0360 | 0.0520 | 0.0400 |
| 3000 | 0.0573 | 0.0480 | 0.0507 | 0.0467 | 0.0440 | 0.0453 | 0.0487 | 0.0447 |
| R | T50 | |||||||
| 500 | 0.9200 | 0.9240 | 0.8640 | 0.8880 | 0.8480 | 0.8160 | 0.7760 | 0.7800 |
| 1000 | 0.9000 | 0.9000 | 0.8860 | 0.8980 | 0.8000 | 0.8320 | 0.7840 | 0.7840 |
| 3000 | 0.9247 | 0.9253 | 0.8713 | 0.8760 | 0.8247 | 0.8273 | 0.7980 | 0.8013 |
| R | T100 | |||||||
| 500 | 0.9920 | 0.9920 | 0.9920 | 0.9880 | 0.9840 | 0.9800 | 0.9880 | 0.9720 |
| 1000 | 0.9880 | 0.9880 | 0.9860 | 0.9840 | 0.9800 | 0.9720 | 0.9740 | 0.9840 |
| 3000 | 0.9893 | 0.9893 | 0.9920 | 0.9827 | 0.9820 | 0.9773 | 0.9767 | 0.9747 |
| R | T500 | |||||||
| 500 | 1.0000 | 1.0000 | 1.0000 | 0.9960 | 1.0000 | 0.9960 | 1.0000 | 1.0000 |
| 1000 | 1.0000 | 1.0000 | 1.0000 | 1.0000 | 0.9980 | 1.0000 | 0.9980 | 1.0000 |
| 3000 | 1.0000 | 0.9987 | 1.0000 | 1.0000 | 0.9993 | 1.0000 | 0.9993 | 0.9993 |
| R | T1000 | |||||||
| 500 | 1 | 1 | 1 | 1 | 1 | 1 | 1 | 1 |
| 1000 | 1 | 1 | 1 | 1 | 1 | 1 | 1 | 1 |
| 3000 | 1 | 1 | 1 | 1 | 1 | 1 | 1 | 1 |
6 Final comments. Reliable inference from our approach
The simulation study illustrates six key aspects of our approach, briefly summarized in points (i)–(vi) below:
The tight property under of the random projection sequence
allows the application of Prokhorov Theorem to prove the convergence, as of to in the space In particular, the asymptotic Gaussian distribution of under follows from this result. This result is illustrated in Section 5.1 from Theorem 1.2.1 in [10].
The crucial role played by the design of the test statistic operator in the derivation of the conditions assumed to obtain consistency of the test (see Proposition 1 and Theorem 2). The simulation study also reveals that the additional conditions assumed in Theorem 3 lead to universal a.s. divergence rates. These rates are not affected by the localization of the dominant eigenspace, or the value of the parameter chosen, under the bandwidth parameter scenario when Table 5 visualizes this fact for parameter values Specifically, in the three examples analyzed, the sample values of projected into are displayed under these three bandwidth parameter scenarios, for functional sample sizes Note that, although, as expected, the sample values of slightly increase when increases (see Examples 1–3), no significant differences are observed in the sample divergence rate of between the three values of parameter analyzed. Furthermore, under condition when one can observe the invariance of the sample divergence rate against the location of the dominant eigenspace. Keeping in mind that, as illustrated in [26], the regular spectral factor which here is represented by the SPHARMA(p,q) functional spectrum, has not impact in the asymptotic analysis, one can conclude the findings in Table 5 are representative.
Table 5: projected into T Example 1 Example 2 Example 3 1.6885(1.0e+04) 1.6384(1.0e+04) 1.6438(1.0e+04) 4.3549(1.0e+07) 4.2251(1.0e+07) 4.2290(1.0e+07) 1.7493(1.0e+08) 1.6944(1.0e+08) 1.6989(1.0e+08) 1.8067 (1.0e+04) 1.7733(1.0e+04) 1.7693(1.0e+04) 4.4789(1.0e+07) 4.3747(1.0e+07) 4.3510(1.0e+07) 1.7984(1.0e+08) 1.7470(1.0e+08) 1.7476(1.0e+08) 2.0296(1.0e+04) 2.0109(1.0e+04) 1.9993(1.0e+04) 4.5133(1.0e+07) 4.4271(1.0e+07) 4.4138(1.0e+07) 1.8040 (1.0e+08) 1.7518(1.0e+08) 1.7624(1.0e+08) -
As expected, when higher orders of singularity are displayed at zero frequency, beyond the restriction a faster divergence of the sample values of the Hilbert–Schmidt operator norm of and of its diagonal projections, is observed.
The testing approach adopted shows good empirical size and power properties for finite functional samples, as reported in Tables 3 and 4. Namely, Table 3, in the simulation study undertaken, shows empirical test sizes very close to the theoretical value for the minimum sample size considered, and for the number of repetitions Table 4 displays, for empirical powers in the interval at the eight random directions tested, and for the number of repetitions Summarizing, as illustrated in Section 5.4, for relatively small functional sample sizes, reliable inference based on our functional spectral nonparametric approach is possible. On the other hand, asymptotic properties like consistency of the test are verified to hold, as given in Table 1, displaying the increasing order of magnitude of the Hilbert–Schmidt operator norm of our test statistics, which is around for the smallest functional sample size in all examples. By the same reasons explained in (ii), i.e., the invariance of the results displayed in Tables 3 and 4 against the location of the dominant eigenspace ( and respectively in Examples 1,2,3), and the absence of asymptotic impact of the choice of the SPHARMA(p,q) functional spectrum, one can conclude the representativeness of the numerical results reflected in Tables 3 and 4.
We remark that all computations involved in the simulation study undertaken, in particular, in the implementation of the proposed inference tools and testing approach, have been achieved in terms of a unique orthonormal basis, given by the eigenfunctions of the Laplace Beltrami operator. Thus, we have worked under the scenario where the eigenfunctions of the elements of the covariance and spectral density operator families are known. This fact constitutes an important advantage of the analyzed setting, avoiding the use of empirical eigenfunction bases. We have also worked under the context of fully observed functional data. The case of sparse discretely observed and contaminated functional data can be addressed from the nonparametric series least–squares estimation of our functional data set, and plug–in implementation of our test statistics (see, e.g., [46] and [37]). Specifically, under suitable restrictions on the local Hölder regularity of our functional data set, and on the supremum norm of the sieve basis elements, as well as on the pure point spectral properties of the autocovariance matrix of the random sieve basis (involved in the nonparametric series least–squares estimation of the discretely observed functional data set), one can derive similar asymptotic results. To this aim, a suitable manifold random uniform sampling design must be considered. The time–varying manifold sampling frequency must display a faster divergence rate than the time–varying sieve basis dimension, but slower than the functional sample size. Under these conditions, similar results on asymptotic bias analysis can be obtained under depending on the almost surely uniform convergence rates to the theoretical values of our functional data, involved in their nonparametric series least–squares estimation.
As commented in the Introduction, the presented approach allows the statistical inference from spatiotemporal data sets embedded into the sphere. Thus, the non–euclidean spatial statistical analysis of such data sets can be performed enhancing geometrical interpretability, avoiding the usual transformations of longitudes and latitudes of the data to work in a cartesian reference coordinate system. Under the invariance properties assumed in our setting, the case of discretely observed functional data can be addressed considering the sieve basis constructed from the eigenfunctions of the Laplace Beltrami operator. This sieve basis allows an easier geometrical implementation (see, e.g., [18], and [46] for alternative sieve bases in nonparametric series least–squares ridge regression in an euclidean spatial setting). An important dimension reduction is obtained in terms of this sieve basis, which is crucial in the reconstruction of high–dimensional data sets. This fact constitutes another remarkable feature of our approach that reduces computational burden, allowing the implementation of resampling techniques in real data applications. Note that temporal information can be incorporated under our functional time series framework, extending recent developments in the purely spatial statistical euclidean context (see, e.g., [18]). Just to mention a motivating data example for implementation of our approach, the authors in [18] analyze a georeferenced spatiotemporal discretely observed population data set to predict nighttime population in Tokyo. They implement series spatial ridge regression estimation after monthly averaging the data, ignoring time information that is crucial in this prediction problem. Our functional spectral nonparametric approach allows time information to be processed in an efficient way in an non–euclidean setting as commented in (iv) and (v).
Acknowledgements. This work has been supported in part by projects PID2022–142900NB-I00 and PID2020-116587GB-I00, financed by MCIU/AEI/10.13039/501100011033 and by FEDER UE, and CEX2020-001105-M MCIN/AEI/10.13039/501100011033), as well as supported by grant ED431C 2021/24 (Grupos competitivos) financed by Xunta de Galicia through European Regional Development Funds (ERDF).
Appendix. Proof of the results
Appendix A Proofs of the results in Section 2
A.1 Proof of Theorem 1
Proof. From Lemma 2.1,
where denotes the convergence in distribution. Here, are jointly zero–mean complex Gaussian elements in with covariance kernel (23).
Let us consider the orthonormal basis of eigenfunctions of the Laplace–Beltrami operator on From equation (23), applying invariance property leading to
we obtain, for
From (48), applying Cauchy–Schwartz inequality, and the orthonormality of the basis of eigenfunctions of the Laplace Beltrami operator, the following inequalities hold: For every and
under SRD, and for certain and with as As before, denotes the weighted periodogram operator with kernel In equation (49), we have considered sufficiently large to apply the identity for and (see Lemma F11 in the Supplementary Material of [28]). Thus, under assuming the conditions in Lemma 2.1, the sequence is tight. Hence, the convergence, as of to the Gaussian random operator with kernel (see equation (45)), in the norm of the space follows from Prokhorov Theorem. Here, denotes the space of zero–mean second–order –valued random variables with the norm
Let us consider
for and where for a bounded linear operator on a separable Hilbert space for every Here, as before, is the Gaussian random element with random kernel introduced in equation (45) for Then, applying Cauchy–Schwartz inequality in the space of complex–valued zero–mean second–order random variables on and Jensen’s inequality, we obtain
under for certain positive constant where the last inequality follows from equations (47) and (48), for sufficiently large.
Applying Dominated Convergence Theorem, we then obtain
in view of the convergence in of to Thus, the convergence in distribution of to holds.
Appendix B Proofs of the results in Section 3
B.1 Proof of Lemma 2
Proof. Let us consider
where the sequence of functions
defines the frequency–varying pure point spectra of the operator family for every with denoting the periodogram operator (see equation (4)), and denoting the expectation under the alternative
Under for every Applying well–known properties of Féjer kernel, we then obtain, as
From Young convolution inequality in for each and
under since Thus, for every and
To apply Dominated Convergence Theorem in (52), the following additional inequalities are considered, obtained from triangle inequality, Young convolution inequality for functions in and Jensen’s inequality, keeping in mind that a.s. in for every
Also, in a similar way,
Under from equation (16),
From equations (53)–(57), one can apply Dominated Convergence Theorem in equation (52), obtaining
The rate of convergence to zero of the bias is now obtained in the time domain. Let be defined as
The function sequence
pointwise converges, as to with rate of convergence and satisfies, for every
B.2 Proof of Corollary 1
Proof.
Applying Lemma 3.1, and Lemmas F10 and F12 of Appendix F in the Supplementary Material of [28],
as we wanted to prove.
B.3 Proof of Lemma 3
Proof.
Under Assumption I, there exists an orthonormal basis of such that (see [17])
Furthermore,
with and otherwise. In (65), we have considered the change of variable and Denote, for every and
where, for satisfies (64). From equations (64)–(65), applying Fourier transform inversion formula, for each
As uniformly in
where denotes the Dirac Delta distribution, defining the kernel of the identity operator on Using the notation
equation (66) can be rewritten as
Note that, for with sufficiently large,
since for every with Under Assumption I, applying Parseval identity,
Hence, from (70)–(71), applying Dominated Convergence Theorem,
and, as
It then follows that, as the norm
goes to zero, with
in the norm of the space where denotes the cumulant spectral density operator of order of under introduced in equation (30).
Appendix C Proofs of the results in Section 4
C.1 Proof of Proposition 1
Proof. Under we have for every From Lemma 3.1, considering sufficiently large,
where and respectively denote the frequency varying eigenvalues of the weighted periodogram operator and the mean operator Here, means that the two sequences and have the same limit as From (75),
where denotes the space of bounded linear operators on
C.2 Proof of Theorem 2
Proof. From Lemmas 3.1 and 3.3, applying trace formula, for sufficiently large,
where, as before, means that the two sequences and have the same limit as
C.3 Proof of Corollary 2
Proof. Applying triangle and Jensen inequalities, we obtain from Corollary 3.2 and Theorem 4.2,
C.4 Proof of Theorem 3
Proof. The proof of this result shares some ideas with the proof of Theorem 2 of [16], formulated in the time domain for real–valued time series. Specifically, the test statistic operator is rewritten as
where means the composition of operators, denotes the identity operator on the space and is the inverse of operator
Our strategy in the proof of this result consists of first proving, under the divergence, in the norm of the space of operator
Then, under the conditions of Theorem 4.2, we derive the convergence to zero, as of random operator
in the space which holds with a suitable rate under and allowing the application of Borell Cantelli Lemma to ensure almost surely convergence. Specifically, from Proposition 4.1, as
For the random operator in (80), the following inequality holds:
From equation (76), as
and, from Theorem 4.2,
For each applying Jensen inequality, in terms of the uniform probability measure on the interval
where denotes expectation under the uniform probability measure on the interval and is a convex function. Thus, from equations (82)–(85),
From equation (86), applying Chebyshev’s inequality,
Since hence, and, for with and From equation (87), Borel–Cantelli lemma then leads to
as The a.s. divergence of as follows from equations (79), (81) and (88).
C.5 Proof of Lemma 4
Proof. For consider a Gaussian random element in the space with kernel satisfying (23). Hence, from (47),
for every Thus, the diagonal coefficients define the eigenvalues of the autocovariance operator (89). Since is a compact set, and is a trace positive semidefinite self–adjoint operator under the orthogonal expansion
holds in the space The random Fourier coefficients are given by